# Entanglement suppression and quantum scars in a three-oscillator gravitational analogue

### v1

@@ Appendix B Solution to the general three Harmonic oscillator model
In this section, we present the explicit solution for the dynamics of the system discussed in main text [53]. The Hamiltonian quadratic form ($H$) of the system is composed of purely real parameters, as a result the entire covariance formalism can be reformulataed assuming all the matrices to be real. Thus the set of Eqs. (24) and Eqs. (24) and (28) reduce to
@@
where $\bm{U}$ and $\bm{V}$ are given by Eq.(4) in the main text [53]. To check whether the system of equations at hand are consistent, we can count the number of independent unknowns and independent equations in them. Due to symmetry of the matrices $\bm{U}$ and $\bm{V}$, Eqs. (31c) - (31d) constitute 12 independent equations in 21 unknowns. However, the 21 unknowns are all not independent and are constrained by Eqs. (31a) – (31b) which are 9 in number. Thus, we have (21 - 9) = 12 unknowns and 12 independent equations.
@@
For the given choice of matrices $\bm{U}$ and $\bm{V}$, we can explicitly find our the symplectic eigenvalues of $H$ by writing the characterisitic polynomial of the matrix $i\Omega H$ under the assumptions $\omega_{1}>\lambda_{1},\omega_{2}>\lambda_{2}$, which gives
$f(\lambda)$ $=$ $(\lambda^{2}+\lambda_{1}^{2}-\omega_{1}^{2})(\lambda^{4}+R(\omega_{1}-\lambda_{1})(\omega_{2}-\lambda_{2})$ (34)
$+\lambda^{2}(\lambda_{1}^{2}+\lambda_{2}^{2}-\omega_{1}^{2}-\omega_{2}^{2}))$
$R$ $=$ $(\omega_{1}+\lambda_{1})(\omega_{2}+\lambda_{2})-4(g_{1}^{2}+g_{2}^{2})$
@@
$k_{1}$ $=\sqrt{\omega_{1}^{2}-\lambda_{1}^{2}}$ (35)
$k_{2}$ $=\frac{1}{\sqrt{2}}\sqrt{(\omega_{1}^{2}+\omega_{2}^{2}-\lambda_{1}^{2}-\lambda_{2}^{2})-\sqrt{Q}}$ (36)
$k_{3}$ $=\frac{1}{\sqrt{2}}\sqrt{(\omega_{1}^{2}+\omega_{2}^{2}-\lambda_{1}^{2}-\lambda_{2}^{2})+\sqrt{Q}}$ (37)
$Q$ $=(\omega_{1}^{2}+\omega_{2}^{2}-\lambda_{1}^{2}-\lambda_{2}^{2})^{2}-$
$4(\omega_{1}-\lambda_{1})(\omega_{2}-\lambda_{2})[(\omega_{1}+\lambda_{1})(\omega_{2}+\lambda_{2})-4(g_{1}^{2}+g_{2}^{2})]$
@@
$g_{1}^{2}+g_{2}^{2}\leq\frac{1}{4}(\omega_{1}+\lambda_{1})(\omega_{2}+\lambda_{2})$ (38)
@@
$\beta_{11}$ $=\frac{\alpha_{11}(k_{1}-\omega_{1}+\lambda_{1})}{(k_{1}+\omega_{1}-\lambda_{1})}$ (39)
$\beta_{12}$ $=\frac{\alpha_{12}(k_{1}-\omega_{2}+\lambda_{2})}{(k_{1}+\omega_{2}-\lambda_{2})}$ (40)
$\beta_{13}$ $=\frac{\alpha_{13}(k_{1}-\omega_{1}+\lambda_{1})}{(k_{1}+\omega_{1}-\lambda_{1})}$ (41)
$\beta_{21}$ $=\frac{\alpha_{21}(k_{2}-\omega_{1}+\lambda_{1})}{(k_{2}+\omega_{1}-\lambda_{1})}$ (42)
$\beta_{22}$ $=\frac{\alpha_{22}(k_{2}-\omega_{2}+\lambda_{2})}{(k_{2}+\omega_{2}-\lambda_{2})}$ (43)
$\beta_{23}$ $=\frac{\alpha_{23}(k_{2}-\omega_{1}+\lambda_{1})}{(k_{2}+\omega_{1}-\lambda_{1})}$ (44)
$\beta_{31}$ $=\frac{\alpha_{31}(k_{3}-\omega_{1}+\lambda_{1})}{(k_{3}+\omega_{1}-\lambda_{1})}$ (45)
$\beta_{32}$ $=\frac{\alpha_{32}(k_{3}-\omega_{2}+\lambda_{2})}{(k_{3}+\omega_{3}-\lambda_{2})}$ (46)
$\beta_{33}$ $=\frac{\alpha_{33}(k_{3}-\omega_{1}+\lambda_{1})}{(k_{3}+\omega_{1}-\lambda_{1})}$ (47)
@@
$(\alpha_{11}^{2}+\alpha_{13}^{2})\frac{4k_{1}(\omega_{1}-\lambda_{1})}{(k_{1}+\omega_{1}-\lambda_{1})^{2}}\;+\alpha_{12}^{2}\frac{4k_{1}(\omega_{2}-\lambda_{2})}{(k_{1}+\omega_{2}-\lambda_{2})^{2}}=1$ (48)
$(\alpha_{21}^{2}+\alpha_{23}^{2})\frac{4k_{2}(\omega_{1}-\lambda_{1})}{(k_{2}+\omega_{1}-\lambda_{1})^{2}}\;+\alpha_{22}^{2}\frac{4k_{2}(\omega_{2}-\lambda_{2})}{(k_{2}+\omega_{2}-\lambda_{2})^{2}}=1$ (49)
$(\alpha_{31}^{2}+\alpha_{33}^{2})\frac{4k_{3}(\omega_{1}-\lambda_{1})}{(k_{3}+\omega_{1}-\lambda_{1})^{2}}\;+\alpha_{32}^{2}\frac{4k_{3}(\omega_{2}-\lambda_{2})}{(k_{3}+\omega_{2}-\lambda_{2})^{2}}=1$ (50)
@@
$\frac{(\alpha_{11}\alpha_{21}+\alpha_{13}\alpha_{23})(\omega_{1}-\lambda_{1})}{(k_{1}+\omega_{1}-\lambda_{1})(k_{2}+\omega_{1}-\lambda_{1})}\;+\frac{\alpha_{12}\alpha_{22}(\omega_{2}-\lambda_{2})}{(k_{1}+\omega_{2}-\lambda_{2})(k_{2}+\omega_{2}-\lambda_{2})}=0$ (51)
$\frac{(\alpha_{11}\alpha_{31}+\alpha_{13}\alpha_{33})(\omega_{1}-\lambda_{1})}{(k_{1}+\omega_{1}-\lambda_{1})(k_{3}+\omega_{1}-\lambda_{1})}\;+\frac{\alpha_{12}\alpha_{32}(\omega_{2}-\lambda_{2})}{(k_{1}+\omega_{2}-\lambda_{2})(k_{3}+\omega_{2}-\lambda_{2})}=0$ (52)
$\frac{(\alpha_{21}\alpha_{31}+\alpha_{23}\alpha_{33})(\omega_{1}-\lambda_{1})}{(k_{2}+\omega_{1}-\lambda_{1})(k_{3}+\omega_{1}-\lambda_{1})}\;+\frac{\alpha_{22}\alpha_{32}(\omega_{2}-\lambda_{2})}{(k_{2}+\omega_{2}-\lambda_{2})(k_{3}+\omega_{2}-\lambda_{2})}=0$ (53)
@@
$\alpha_{1j}$ $\dfrac{(k_{1}+\omega_{1}-\lambda_{1})x_{1}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$ 0 $-\dfrac{(k_{1}+\omega_{1}-\lambda_{1})y_{1}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$
$\alpha_{2j}$ $\dfrac{(k_{2}+\omega_{1}-\lambda_{1})y_{1}y_{2}}{2\sqrt{k_{2}(\omega_{1}-\lambda_{1})}}$ $\dfrac{(k_{2}+\omega_{2}-\lambda_{2})x_{2}}{2\sqrt{k_{2}(\omega_{2}-\lambda_{2})}}$ $\dfrac{(k_{2}+\omega_{1}-\lambda_{1})x_{1}y_{2}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$
$\alpha_{3j}$ $-\dfrac{(k_{3}+\omega_{1}-\lambda_{1})y_{1}x_{2}}{2\sqrt{k_{3}(\omega_{1}-\lambda_{1})}}$ $\dfrac{(k_{2}+\omega_{2}-\lambda_{2})y_{2}}{2\sqrt{k_{3}(\omega_{2}-\lambda_{2})}}$ $-\dfrac{(k_{3}+\omega_{1}-\lambda_{1})x_{1}x_{2}}{2\sqrt{k_{3}(\omega_{1}-\lambda_{1})}}$
$\beta_{1j}$ $\dfrac{(k_{1}-\omega_{1}+\lambda_{1})x_{1}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$ 0 $-\dfrac{(k_{1}-\omega_{1}+\lambda_{1})y_{1}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$
$\beta_{2j}$ $\dfrac{(k_{2}-\omega_{1}+\lambda_{1})y_{1}y_{2}}{2\sqrt{k_{2}(\omega_{1}-\lambda_{1})}}$ $\dfrac{(k_{2}-\omega_{2}+\lambda_{2})x_{2}}{2\sqrt{k_{2}(\omega_{2}-\lambda_{2})}}$ $\dfrac{(k_{2}-\omega_{1}+\lambda_{1})x_{1}y_{2}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$
$\beta_{3j}$ $-\dfrac{(k_{3}-\omega_{1}+\lambda_{1})y_{1}x_{2}}{2\sqrt{k_{3}(\omega_{1}-\lambda_{1})}}$ $\dfrac{(k_{2}-\omega_{2}+\lambda_{2})y_{2}}{2\sqrt{k_{3}(\omega_{2}-\lambda_{2})}}$ $-\dfrac{(k_{3}-\omega_{1}+\lambda_{1})x_{1}x_{2}}{2\sqrt{k_{3}(\omega_{1}-\lambda_{1})}}$
@@
$x_{1}$ $=\cos{\theta_{1}}=\frac{g_{1}^{2}}{\sqrt{g_{1}^{2}+g_{2}^{2}}}$ (54)
$y_{1}$ $=\sin{\theta_{1}}=\frac{g_{2}^{2}}{\sqrt{g_{1}^{2}+g_{2}^{2}}}$ (55)
$x_{2}$ $=\cos{\theta_{2}}=\frac{(k_{2}^{2}+\lambda_{1}^{2}-\omega_{1}^{2})}{\sqrt{T}}$ (56)
$y_{2}$ $=\sin{\theta_{2}}=\frac{2\sqrt{(\omega_{1}-\lambda_{1})(\omega_{2}-\lambda_{2})(g_{1}^{2}+g_{2}^{2})}}{\sqrt{T}}$ (57)
$T$ $=(k_{2}^{2}+\lambda_{1}^{2}-\omega_{1}^{2})^{2}+4(\omega_{1}-\lambda_{1})(\omega_{2}-\lambda_{2})(g_{1}^{2}+g_{2}^{2})$
@@ Appendix C Quantum Measures
For Gaussian states, entanglement is evaluated via the partial transpose operation, implemented by applying the matrix $\Lambda$ [54], yielding the partially transposed covariance matrix:
@@
The logarithmic negativity is computed from the symplectic eigenvalues $\tilde{\nu}_{i}(t)$ of $\tilde{\bm{\sigma}}_{A-C}(t)$, which are the eigenvalues of $i\Omega^{\prime}\tilde{\bm{\sigma}}_{A-C}(t)$, where $\Omega^{\prime}$ is the reduced symplectic form [24]. Explicitly,
@@
and $\bm{\sigma}_{0}$ corresponds to the covariance matrix of a single-mode Gaussian measurement on subsystem C. The minimum value of the determinant of $\epsilon$ is given in Eq. (67), where $\mathcal{A}=\det(\bm{\sigma}_{A})$, $\mathcal{B}=\det(\bm{\sigma}_{B})$, $\mathcal{C}=\det(\bm{\sigma}_{C})$, and $\mathcal{D}=\det(\bm{\sigma}_{A-C})$. The fidelity between two covariance matrices $V_{1}$ and $V_{2}$ is defined as [31]:

### v3

@@ Appendix B Solution to the general three Harmonic oscillator model
In this section, we present the explicit solution for the dynamics of the system discussed in main text. The Hamiltonian quadratic form ($H$) of the system is composed of purely real parameters, as a result the entire covariance formalism can be reformulataed assuming all the matrices to be real. Thus the set of Eqs. (26) and Eqs. (26) and (30) reduce to
@@
where $\bm{U}$ and $\bm{V}$ are given by Eq. (4). To check whether the system of equations at hand are consistent, we can count the number of independent unknowns and independent equations in them. Due to symmetry of the matrices $\bm{U}$ and $\bm{V}$, Eqs. (33c) - (33d) constitute 12 independent equations in 21 unknowns. However, the 21 unknowns are all not independent and are constrained by Eqs. (33a) – (33b) which are 9 in number. Thus, we have (21 - 9) = 12 unknowns and 12 independent equations.
@@
For the given choice of matrices $\bm{U}$ and $\bm{V}$, we can explicitly find our the symplectic eigenvalues of $H$ by writing the characterisitic polynomial of the matrix $i\Omega H$ under the assumptions $\omega_{A}>\lambda_{A},\omega_{B}>\lambda_{B}$, which gives
$f(\lambda)$ $=$ $(\lambda^{2}+\lambda_{A}^{2}-\omega_{A}^{2})(\lambda^{4}+R(\omega_{A}-\lambda_{A})(\omega_{B}-\lambda_{B})$ (36)
$+\lambda^{2}(\lambda_{A}^{2}+\lambda_{B}^{2}-\omega_{A}^{2}-\omega_{B}^{2}))$
$R$ $=$ $(\omega_{A}+\lambda_{A})(\omega_{B}+\lambda_{B})-4(g_{A}^{2}+g_{B}^{2})$
@@
$k_{1}$ $=\sqrt{\omega_{A}^{2}-\lambda_{A}^{2}}$ (37)
$k_{2}$ $=\frac{1}{\sqrt{2}}\sqrt{(\omega_{A}^{2}+\omega_{B}^{2}-\lambda_{A}^{2}-\lambda_{B}^{2})-\sqrt{Q}}$ (38)
$k_{3}$ $=\frac{1}{\sqrt{2}}\sqrt{(\omega_{A}^{2}+\omega_{B}^{2}-\lambda_{A}^{2}-\lambda_{B}^{2})+\sqrt{Q}}$ (39)
$Q$ $=(\omega_{A}^{2}+\omega_{B}^{2}-\lambda_{A}^{2}-\lambda_{B}^{2})^{2}$-
$4(\omega_{A}-\lambda_{A})(\omega_{B}-\lambda_{B})[(\omega_{A}+\lambda_{A})(\omega_{B}+\lambda_{B})-4(g_{A}^{2}+g_{B}^{2})]$
@@
$g_{A}^{2}+g_{B}^{2}\leq\frac{1}{4}(\omega_{A}+\lambda_{A})(\omega_{B}+\lambda_{B})$ (40)
@@
$\beta_{11}$ $=\frac{\alpha_{11}(k_{1}-\omega_{A}+\lambda_{A})}{(k_{1}+\omega_{A}-\lambda_{A})}$ (41)
$\beta_{12}$ $=\frac{\alpha_{12}(k_{1}-\omega_{B}+\lambda_{B})}{(k_{1}+\omega_{B}-\lambda_{B})}$ (42)
$\beta_{13}$ $=\frac{\alpha_{13}(k_{1}-\omega_{A}+\lambda_{A})}{(k_{1}+\omega_{A}-\lambda_{A})}$ (43)
$\beta_{21}$ $=\frac{\alpha_{21}(k_{2}-\omega_{A}+\lambda_{A})}{(k_{2}+\omega_{A}-\lambda_{A})}$ (44)
$\beta_{22}$ $=\frac{\alpha_{22}(k_{2}-\omega_{B}+\lambda_{B})}{(k_{2}+\omega_{B}-\lambda_{B})}$ (45)
$\beta_{23}$ $=\frac{\alpha_{23}(k_{2}-\omega_{A}+\lambda_{A})}{(k_{2}+\omega_{A}-\lambda_{A})}$ (46)
$\beta_{31}$ $=\frac{\alpha_{31}(k_{3}-\omega_{A}+\lambda_{A})}{(k_{3}+\omega_{A}-\lambda_{A})}$ (47)
$\beta_{32}$ $=\frac{\alpha_{32}(k_{3}-\omega_{B}+\lambda_{B})}{(k_{3}+\omega_{B}-\lambda_{B})}$ (48)
$\beta_{33}$ $=\frac{\alpha_{33}(k_{3}-\omega_{A}+\lambda_{A})}{(k_{3}+\omega_{A}-\lambda_{A})}$ (49)
@@
$(\alpha_{11}^{2}+\alpha_{13}^{2})\frac{4k_{1}(\omega_{A}-\lambda_{A})}{(k_{1}+\omega_{A}-\lambda_{A})^{2}}\;+\alpha_{12}^{2}\frac{4k_{1}(\omega_{B}-\lambda_{B})}{(k_{1}+\omega_{B}-\lambda_{B})^{2}}=1$ (50)
$(\alpha_{21}^{2}+\alpha_{23}^{2})\frac{4k_{2}(\omega_{A}-\lambda_{A})}{(k_{2}+\omega_{A}-\lambda_{A})^{2}}\;+\alpha_{22}^{2}\frac{4k_{2}(\omega_{B}-\lambda_{B})}{(k_{2}+\omega_{B}-\lambda_{B})^{2}}=1$ (51)
$(\alpha_{31}^{2}+\alpha_{33}^{2})\frac{4k_{3}(\omega_{A}-\lambda_{A})}{(k_{3}+\omega_{A}-\lambda_{A})^{2}}\;+\alpha_{32}^{2}\frac{4k_{3}(\omega_{B}-\lambda_{B})}{(k_{3}+\omega_{B}-\lambda_{B})^{2}}=1$ (52)
@@
$\frac{(\alpha_{11}\alpha_{21}+\alpha_{13}\alpha_{23})(\omega_{A}-\lambda_{A})}{(k_{1}+\omega_{A}-\lambda_{A})(k_{2}+\omega_{A}-\lambda_{A})}\;+\frac{\alpha_{12}\alpha_{22}(\omega_{B}-\lambda_{B})}{(k_{1}+\omega_{B}-\lambda_{B})(k_{2}+\omega_{B}-\lambda_{B})}=0$ (53)
$\frac{(\alpha_{11}\alpha_{31}+\alpha_{13}\alpha_{33})(\omega_{A}-\lambda_{A})}{(k_{1}+\omega_{A}-\lambda_{A})(k_{3}+\omega_{A}-\lambda_{A})}\;+\frac{\alpha_{12}\alpha_{32}(\omega_{B}-\lambda_{B})}{(k_{1}+\omega_{B}-\lambda_{B})(k_{3}+\omega_{B}-\lambda_{B})}=0$ (54)
$\frac{(\alpha_{21}\alpha_{31}+\alpha_{23}\alpha_{33})(\omega_{A}-\lambda_{A})}{(k_{2}+\omega_{A}-\lambda_{A})(k_{3}+\omega_{A}-\lambda_{A})}\;+\frac{\alpha_{22}\alpha_{32}(\omega_{B}-\lambda_{B})}{(k_{2}+\omega_{B}-\lambda_{B})(k_{3}+\omega_{B}-\lambda_{B})}=0$ (55)
@@
$\alpha_{1j}$ $\dfrac{(k_{1}+\omega_{1}-\lambda_{1})x_{1}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$ 0 $\dfrac{(k_{1}+\omega_{1}-\lambda_{1})y_{1}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$
$\alpha_{2j}$ $\dfrac{(k_{2}+\omega_{1}-\lambda_{1})y_{1}y_{2}}{2\sqrt{k_{2}(\omega_{1}-\lambda_{1})}}$ $\dfrac{(k_{2}+\omega_{2}-\lambda_{2})x_{2}}{2\sqrt{k_{2}(\omega_{2}-\lambda_{2})}}$ $\dfrac{(k_{2}+\omega_{1}-\lambda_{1})x_{1}y_{2}}{2\sqrt{k_{2}(\omega_{1}-\lambda_{1})}}$
$\alpha_{3j}$ $\dfrac{(k_{3}+\omega_{1}-\lambda_{1})y_{1}x_{2}}{2\sqrt{k_{3}(\omega_{1}-\lambda_{1})}}$ $\dfrac{(k_{2}+\omega_{2}-\lambda_{2})y_{2}}{2\sqrt{k_{3}(\omega_{2}-\lambda_{2})}}$ $\dfrac{(k_{3}+\omega_{1}-\lambda_{1})x_{1}x_{2}}{2\sqrt{k_{3}(\omega_{1}-\lambda_{1})}}$
$\beta_{1j}$ $\dfrac{(k_{1}-\omega_{1}+\lambda_{1})x_{1}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$ 0 $\dfrac{(k_{1}-\omega_{1}+\lambda_{1})y_{1}}{2\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$
$\beta_{2j}$ $\dfrac{(k_{2}-\omega_{1}+\lambda_{1})y_{1}y_{2}}{2\sqrt{k_{2}(\omega_{1}-\lambda_{1})}}$ $\dfrac{(k_{2}-\omega_{2}+\lambda_{2})x_{2}}{2\sqrt{k_{2}(\omega_{2}-\lambda_{2})}}$ $\dfrac{(k_{2}-\omega_{1}+\lambda_{1})x_{1}y_{2}}{2\sqrt{k_{2}(\omega_{1}-\lambda_{1})}}$
$\beta_{3j}$ $\dfrac{(k_{3}-\omega_{1}+\lambda_{1})y_{1}x_{2}}{2\sqrt{k_{3}(\omega_{1}-\lambda_{1})}}$ $\dfrac{(k_{2}-\omega_{2}+\lambda_{2})y_{2}}{2\sqrt{k_{3}(\omega_{2}-\lambda_{2})}}$ $\dfrac{(k_{3}-\omega_{1}+\lambda_{1})x_{1}x_{2}}{2\sqrt{k_{3}(\omega_{1}-\lambda_{1})}}$
@@
[…]
$T$ $=(k_{2}^{2}+\lambda_{A}^{2}-\omega_{A}^{2})^{2}+4(\omega_{A}-\lambda_{A})(\omega_{B}-\lambda_{B})(g_{A}^{2}+g_{B}^{2})$
@@ Appendix C Quantum Measures
For Gaussian states, entanglement is evaluated via the partial transpose operation, implemented by applying the matrix $\Lambda$ Simon [2000], yielding the partially transposed covariance matrix:
@@
The logarithmic negativity is computed from the symplectic eigenvalues $\tilde{\nu}_{i}(t)$ of $\tilde{\bm{\sigma}}_{A-C}(t)$, which are the eigenvalues of $i\Omega^{\prime}\tilde{\bm{\sigma}}_{A-C}(t)$, where $\Omega^{\prime}$ is the reduced symplectic form Vidal and Werner [2002], Plenio [2005]. Explicitly,
@@
and $\bm{\sigma}_{0}$ corresponds to the covariance matrix of a single-mode Gaussian measurement on subsystem C. The minimum value of the determinant of $\epsilon$ is given in Eq. (69), where $\mathcal{A}=\det(\bm{\sigma}_{A})$, $\mathcal{B}=\det(\bm{\sigma}_{B})$, $\mathcal{C}=\det(\bm{\sigma}_{C})$, and $\mathcal{D}=\det(\bm{\sigma}_{A-C})$. The fidelity between two covariance matrices $V_{1}$ and $V_{2}$ is defined as Banchi et al. [2015]: